OpenStudy (anonymous):

write and explicit definition for the sequence -4, 1, 6, 11 a. -4+5(n+1) b.-4+5(n-1) c.-4+(n-1) d.-4+5n

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle (1)-(-4)= (6)-(1)= (11)-(11)=5 }\) so you can for the least part that it is a geometric or arithmetic sequence ?